4. Solve the given system of equations. (10 points each) dx dy 6- + dt + 3y = 0 dt dx dt + 4x - y = 0
Answer #1


eigen vector corresponding to x=-12 v=-su (3))-3) 8u+v=o 244 +3 = 0 is an eigenvector of A for the eigenvalue x=-12 general s